### v1

@@ 3.1.1 Comparing base models
3.1.2 LME vs. reasoning paradigm
Since 2024, it has emerged that the performance of language models does not only scale with training-time compute (i.e. larger models and more training data) but also test-time compute. Allowing an LLM more time to generate responses, break down tasks into multiple steps, and reflect on its reasoning process significantly enhances its performance on reasoning tasks. We test whether this reasoning paradigm is also useful for our specific task: Can an LLM accurately reason about its internal knowledge of regions and firms, perform calculations based on this knowledge, and ultimately provide precise estimates of relevant statistics?
We use the open-source reasoning model Qwen QwQ-32B (Qwen Team, 2025). It is based on Qwen 2.5 32B (Yang et al., 2024), an LLM not tuned to reason using test-time compute. By comparing these two models, we can directly measure the effect of the reasoning paradigm on the accuracy. [^5] In contrast to the previous experiments where we use a completion prompt, we use the question-answering prompt (see Section B.1) for the two Qwen models. We pit the text output of the two models against LME trained on the embeddings of the final token of the prompt of the Qwen base model (Qwen 2.5 32B). Due to the high computational costs of the reasoning model, we only test it on a subsample of 500 entities per variable.
A closer look at the reasoning process reveals that the LLM tries to consider different information to provide the best estimates based on its knowledge base. [^6] For example, when asked for the proportion of delinquent mortgages in Atlantic County, New Jersey in 2019 one of many points the reasoning model made was “Wait, I think that in 2019, the national delinquency rate was around 3-4%, but that’s the national average. Atlantic County might be different, especially if it’s a coastal area that might have been affected by Hurricane Sandy in 2012, but that’s a few years prior. Maybe the recovery from that could affect 2019 numbers? Not sure.” Another sentence from the same answer: “Hmm, I’m not entirely sure, but I think the national average was around 1.8%, and New Jersey might be a bit higher. Since Atlantic County is a county with some economic challenges, maybe it’s a bit higher than the state average.”. However, despite the efforts, we do not observe a consistent increase in performance.
Figure 4 shows the results for our two main datasets. Overall, we do not see a consistent improvement in performance when using the reasoning model to generate the text output. However, the reasoning model performs substantially better on the three US county variables on which the non-reasoning model performs worst. LME beats the text output of both models on most variables. The difference in computational costs is significant. The reasoning model uses more tokens to answer the question, a median of 1767 output tokens, whereas Qwen 2.5 answers the question with median of 8 tokens.
@@ 3.1.3 Learning curves
Figure 6: Comparison between text output and baseline LME with different training samples
The cross-validation results show that we can infer statistical variables more accurately from the embeddings than from the text output. However, the 5-fold cross-validation setting assumes that 80% of the values on the variable we are predicting are available to the modeller. If LME can reach a high accuracy with only a small number of labelled data points, its relevance for data processing tasks will be much higher. To test this we conduct a learning curve and estimate the performance as a function of the number of training samples. Figure 5 compares the text output (red line) against the LME when trained on different numbers of observations in training samples for some variables in the US counties and US firm datasets. LME often reaches its maximum performance with only a small number of training samples. Dimensionality reduction via PCA only helps the performance on very small sample sizes and when using only two components. Figure 6 shows the learning curve results for all datasets comparing the text output to baseline LME at a training sample size of 25, 50, and 100 observations. With only 50 samples, LME is superior on nearly all of the variables.
@@ 3.2.1 Learning from embeddings of other variables
The previous analyses assume that we have at least some labelled instances of a variable $v$ to estimate its values on other instances using the embeddings. Here, we test whether we can reliably estimate a regional or firm variable without access to any values of this variable. If the embeddings encode numbers in a consistent way across different variables, we should be able to predict variable $v$ using a model learned from other variables. More formally, given all variables $\mathbb{V}$ of a dataset, we train a regression model on embeddings and outcome values of variables $\mathbb{V}\setminus v$ and test the models on variable $y_{v}$. Training observations are described as $\{(y_{v},E_{v}):v\in\mathbb{V}\setminus v_{test}\}$ and test observations as $\{(y_{v},E_{v}):v=v_{test}\}$. Note that $E_{v}\neq E_{w},\forall v\neq w$, as we use the completion prompt to obtain the embeddings, which also contains the name of the variable of interest $v$.
To ensure our model is flexible enough to learn generalised embeddings, we train a neural network on the embeddings. We chose a neural network with two layers (128 neurons and 32 neurons, Leaky ReLU activation function) as our transfer learning model. We use the Adam optimiser with an initial learning rate of $10^{-5}$. To avoid overfitting, we only learn for 20 epochs and apply a strong dropout of 0.5 on the first layer. We test this approach on all five datasets. Here, and in the following transfer learning analyses, all variables are transformed as described in Section 2.1 before they are normalised (mean of zero, standard deviation of one).
The left panel of Figure 7 compares the performance (Spearman correlation) of this approach to the text prediction. Each dot reflects one variable, for which a separate model was trained on all other variables of the same dataset. This approach to transfer learning does not beat the text prediction. While transfer learning is superior in some cases, it falls behind the text output most of the time. We obtain similar results, when using a linear model instead of the neural network. We also test another simple transfer learning approach, where we predict variable $y$ in Dataset $A$ from the same variable of different datasets. For example, we train a model on the unemployment rate of the UK and Europe to predict the unemployment rate in the US. However, this transfer learning approach is also inferior to the text prediction on average, and sometimes by a large margin.
@@ 3.2.2 Learning from text output
Given the poor performance of the simple transfer learning approaches, we refine our method to exploit information about the variable we predict ($v_{test}$), even without access to ground truth values. Concretely, we use the text outputs of the LLM as noisy or biased labels of the variable of interest in the training set ($\overset{txt}{y_{test}}$). For the remaining variables we use the ground truth labels, as above. Formally, let the training set be $\{([y_{v},\overset{txt}{y}_{v_{test}}],[E_{v},E_{v_{test}}]):v\in\mathbb{V}\setminus v_{test}\}$ and the test set be $\{(y_{v},E_{v}):v=v_{test}\}$. Even though the text labels are often inaccurate, they might provide enough information for the model to learn from the embeddings of the specific prompt of $v_{test}$.
We use the same neural network set-up as above. The right panel of Figure 7 compares this performance of this transfer learning approach to the text output. This transfer learning approach almost always outperforms the text prediction even when the text prediction itself is accurate. On average, transfer learning beats the text output by 7.2 percentage points. This is an important finding. We can extract the knowledge about a specific economic or financial variable from LLM embeddings without any labelled instances of that variable. The result is robust to the choice of the hyperparameters of the neural network and we observe qualitatively similar, but slightly inferior performance, when using a linear model instead of the neural network.
@@ 4.1 Imputation
Our analyses have established that embeddings contain rich knowledge about regions and firms. Here we test whether this generalised knowledge can be exploited to improve imputation accuracy. We randomly mask values of variables in our dataset and use the LLM embeddings to inform standard imputation methods. To obtain a general numeric representation of entities, we use generic prompts (see Section 2.2), which only contain the name of the entity (e.g. region or firm) and the year.
To test whether embeddings help the accuracy of off-the-shelf imputation algorithms, we compare two imputation strategies. The baseline strategy estimates imputed values using only the $K$ numeric features (e.g. population and unemployment rate) in the dataset. The embedding approach extends the feature matrix by appending the first $c$ PCA components of the embeddings.
Our approach to introduce missing values is as follows: Given a dataset with $N$ observations and $K$ variables, all variables are transformed as described in Section 2.1 before they are normalised (mean of zero, standard deviation of 1). First, we sub-sample $k$ features, for each of which we randomly mask $p$of its values, creating the incomplete data $D_{miss}$. We then impute the missing values in $D_{miss}$ using imputation algorithm $m$.
We test three imputation methods $m$: using a low-rank Gaussian copula model (Zhao and Udell, 2020) and iterative multivariate imputation using (i) Bayesian Ridge regression and (ii) random forest (Wilson, 2020). We assess the accuracy of the imputation with the mean absolute difference between ground truth and reconstructed values.
For each imputation method, we conduct a large array of experiments to ensure robustness of our results. Specifically, we consider all possible combinations of the parameters settings $k\in\{1,2,5\},p\in\{0.1,0.25,0.5\}$. To obtain stable results, we replicate the experiments for each parameter combination 25 times, randomly choosing the instances and features which are imputed. We use the first $c=25$ PCA components of the 25th layer of LLama 3 8B-Instruct.
Figure 8 presents the results across all five datasets and parameter settings $m$, $k$, and $p$. Overall, the embedding strategy outperforms the baseline approach, which relies solely on the numeric features. The improvement is particularly notable in some datasets. For instance, in the US counties dataset, the embedding strategy reduces the mean error by 14% on average across parameter settings.
@@ 4.2 Super-resolution
We define a super-resolution problem as the estimation of a statistic at a finer-grained geographical level where actual statistics are not reported or are expected to be imprecise, given values of that statistic at a coarser level. Similarly to imputation, this task has high practical relevance. We can think of many key economic variables that are only published at a higher regional level but that have systematic lower-level regional variation that we want to estimate. LME, exploiting an LLM’s generalised knowledge about regions, could provide accurate estimates in this situation. Unfortunately, we cannot test super-resolution in those scenarios where lower geographical level data do not exist because we require these data to assess the accuracy of our approach.
We approach the super-resolution task by training an LME on a higher regional level (e.g. unemployment rate of US states) to obtain estimates at the lower regional level (e.g. unemployment rate of US counties). We use the completion prompt for both geographical levels.
We test this super-resolution approach on three of our datasets (US, EU, Germany) for which we were able to find data on two regional levels, as shown in Table 3.
@@
We pit this super-resolution approach against the text output for the lower level and against a naive method, which projects the higher region values onto the lower level. For example, the unemployment rate of all counties in Alabama would be set to the unemployment rate reported for the state Alabama. Figure 9 presents the results, ordering the variables by decreasing text performance.
The super-resolution approach performs well, surpassing the text prediction on most variables. A notable outlier is population, where the super-resolution performs worse than random. This result is unique to models in the Llama family, whereas results from Microsoft’s Phi models indicate that super-resolution outperforms the naive approach (Figure AVIII). The naive approach performs well on many of the variables in Germany but falls behind the other two approaches on the US data. Given the small training sample size, in particular for the US and Germany (51 and 38 observations at the higher geographical level, respectively), the performance of the super-resolution approach is impressive. Super-resolution is a particular transfer learning approach and thus supports our previous finding (Section 3.2) that the embeddings can be used for accurate estimation outside the domain of the training set.
@@ 5 Conclusion
We have shown that LLMs know more than they say. A linear model trained on the LLM’s hidden states and only a few dozen labelled samples often outperforms the LLM’s text output in the estimation of statistics – in particular when the target variable is less common statistic. A key contribution of our work is the transfer learning approach which shows that we can estimate target variables more accurately than the text output without any labelled data on that target variable. Finally, we provide robust evidence that the hidden states can be used for data processing tasks. In particular, adding a low-rank representation of the hidden states to the feature matrix consistently improves the performance of standard imputation approaches in all of our datasets. Our super-resolution results show that, in most cases, we can beat the text output in estimating statistics on a lower geographical level by learning from the higher geographical level.
We considered several LLMs with 1 to 70 billion parameters and show that our result that LLMs know more than they say holds across all of them, including a reasoning model that uses extensive test-time compute. We leave it for future work to test whether this finding also holds for the largest and most accurate open source models such LLama 3 405B or DeepkSeek V3 with 671 billion parameters. In this study we exclusively worked with public data sources; it would be interesting to examine the extent to which our results hold for proprietary datasets, which are less likely to be observed in the training corpus of open-source LLMs. Finally, future research could investigate how useful the hidden states representing firms or geographic entities can be for other data processing tasks, such as outlier detection.
@@ A.1 US counties
The data on US counties for the year 2019 is one of our two main datasets and we used a variety of sources to assemble it. We obtained income and GDP data from the website of the Bureau of Economic analysis [^7] https://www.bea.gov/data. The unemployment data was sourced from the Economic Research Service of the US Department of Agriculture [^8] https://www.ers.usda.gov/data-products/county-level-data-sets. Life expectancy data is provided by the Institute for Health Metrics and Evaluation in the Global Health Data Exchange database. We obtained mortgage delinquency rates from the Consumer Financial Protection Bureau. [^9] https://www.consumerfinance.gov/data-research/mortgage-performance-trends/download-the-data Data on the number of businesses was downloaded from the County Business Patterns Tables from the United States Census Bureau. [^10] https://www.census.gov/programs-surveys/cbp/data/tables.html Information on the median rent is available from the website of the Office of Policy Development and research. [^11] https://www.huduser.gov/portal/datasets/50per.html The remaining variables we use come from the R package usdata (Çetinkaya Rundel et al., 2021) which lists the US Census and the American Community Survey as sources.

### v2

@@ 3.2.1 Comparing base models
3.2.2 LME vs. a reasoning model
Since 2024, it has emerged that the performance of language models does not only scale with training-time compute (i.e. larger models and more training data) but also test-time compute. Allowing an LLM more time to generate responses, break down tasks into multiple steps, and reflect on its reasoning process significantly enhances its performance on reasoning tasks. We test whether this reasoning paradigm is also useful for our specific task: Can an LLM accurately reason about its internal knowledge of regions and firms, integrate this information, and ultimately provide precise estimates of relevant statistics?
We use the open-source reasoning model Qwen QwQ-32B (qwq32b). It is based on Qwen 2.5 32B (qwen2.5), an LLM not tuned to reason using test-time compute. By comparing these two models, we can directly measure the effect of the reasoning paradigm on accuracy. [^8] In contrast to the previous experiments where we use a completion prompt, we use the question-answering prompt (see Section B.2) for the two Qwen models in order to elicit reasoning behaviour. We pit the text output of the two models against LME trained on the embeddings of the final token of the prompt of the Qwen base model (Qwen 2.5 32B). Due to the high computational costs of the reasoning model, we only test it on a subsample of 500 entities per variable.
A closer look at the reasoning process reveals that the LLM tries to reconcile different information to provide the best estimates based on its knowledge. [^9] For example, when asked for the proportion of delinquent mortgages in Atlantic County, New Jersey in 2019 one of many points the reasoning model made was “Wait, I think that in 2019, the national delinquency rate was around 3-4%, but that’s the national average. Atlantic County might be different, especially if it’s a coastal area that might have been affected by Hurricane Sandy in 2012, but that’s a few years prior. Maybe the recovery from that could affect 2019 numbers? Not sure.” Another sentence from the same answer: “Hmm, I’m not entirely sure, but I think the national average was around 1.8%, and New Jersey might be a bit higher. Since Atlantic County is a county with some economic challenges, maybe it’s a bit higher than the state average.” However, despite the efforts, we do not observe a consistent increase in performance.
Figure 4 shows the results for our two main datasets. Overall, we do not see a consistent improvement in performance when using the reasoning model to generate the text output. However, the reasoning model performs substantially better on the three US county variables on which the non-reasoning model performs worst. LME beats the text output of both models on most variables. The difference in computational costs is significant. The reasoning model uses a median of 1767 output tokens, whereas Qwen 2.5 answers the question with a median of 8 tokens.
@@ 3.3 Learning curves
Figure 6: Comparison between text output and baseline LME with different training samples.
The cross-validation results show that we can infer statistical variables more accurately from the embeddings than from the text output. However, the 5-fold cross-validation setting assumes that 80% of the values on the variable we are predicting are available to the modeller. If LME can reach a high accuracy with only a small number of labelled data points, its relevance for data processing tasks will be much higher. To test this, we conduct a learning curve analysis, estimating model performance as a function of the number of training samples. Figure 5 compares the text output (red line) with LME trained on progressively larger training sets from the US counties and US firms datasets. LME’s performance often converges with only a small number of training samples. Dimensionality reduction using PCA improves performance on some variables for very small sample sizes and when restricted to two components. Figure 6 summarises the learning-curve results across all datasets, comparing the text output with the baseline LME at training sample sizes of 25, 50, and 100 observations. With only 25 samples, LME outperforms the text output for most variables, showing a mean Spearman correlation 5 percentage points higher.
@@ 3.4.1 Learning from embeddings of other variables
The previous analyses assume that we have at least some labelled instances of a variable to estimate its values on other instances using the embeddings. Here, we test whether we can reliably estimate a regional or firm variable without access to any values of this variable. If the embeddings encode numbers in a consistent way across different variables, we should be able to predict variable $v_{test}$ using a model learned from other variables. More formally, given all variables $\mathbb{V}$ of a dataset with $n$ entities, we train a regression model on embeddings $E_{i}$ and outcome values $y_{i}$ of entities $i$, and test the models on variable $v_{test}$. Training observations are described as $\{(y_{vi},E_{vi}):v\in\mathbb{V}\setminus\{v_{test}\},i=1,...,n\}$ and test observations as $\{(y_{vi},E_{vi}):v=v_{test},i=1,...,n\}$. Note that $E_{v}\neq E_{w},\forall v\neq w$, as we use the completion prompt to obtain the embeddings, which also contains the name of the variables.
To ensure our model is flexible enough to learn generalised embeddings, we train a neural network with two layers (128 neurons and 32 neurons, Leaky ReLU activation function) as our transfer learning model. We use the Adam optimiser with an initial learning rate of $10^{-5}$. To avoid overfitting, we only learn for 20 epochs and apply a strong dropout of 0.5 on the first layer. We test this approach on all five datasets. All variables are transformed as described in Section 2.1 before they are normalised (mean of zero, standard deviation of one).
The left panel of Figure 7 compares the performance (Spearman correlation) of this approach to the text output. Each dot reflects one variable, for which a separate model was trained on all other variables of the same dataset. This approach to transfer learning does not beat the text output. While transfer learning is superior in some cases, it falls behind the text output most of the time. We obtain similar results when using a linear model instead of the neural network. We also test another simple transfer learning approach, where we predict variable $y$ in Dataset $A$ from the same variable of different datasets. For example, we train a model on the unemployment rate of the UK and Europe to predict the unemployment rate in the US. However, this transfer learning approach is also inferior to the text output on average, and sometimes by a large margin.
@@ 3.4.2 Learning from text output
Given the poor performance of the simple transfer learning approaches, we refine our method to exploit information about the variable we predict ($v_{test}$), even without access to ground truth values. Concretely, we use the text outputs of the LLM as noisy or biased labels of the variable of interest in the training set ($\overset{txt}{y}_{v_{test}}$). For the remaining variables we use the ground truth labels, as above. Formally, let $y^{\dagger}_{vi}\;:=\;\begin{cases}y_{vi},&v\neq v_{test}\\ \overset{txt}{y}_{vi},&v=v_{test}\end{cases}$ and the training set be $\{\,(y^{\dagger}_{vi}\,E_{vi})\;:\;v\in\mathbb{V},\ i=1,\ldots,n_{v}\,\}$. The test set, as before, is denoted as $\{(y_{vi},E_{vi}):v=v_{test},i=1,...,n\}$. Even though the text labels are often inaccurate, they might provide enough information for the model to learn from the embeddings of the specific prompt of $v_{test}$.
We use the same neural network set-up as above. The right panel of Figure 7 compares this performance of this transfer learning approach to the text output. This transfer learning approach almost always outperforms the text output even when the text output itself is accurate, on average by 7.2 percentage points. This is an important finding. We can extract knowledge about a specific economic or financial variable from LLM embeddings without any labelled instances of that variable. The result is robust to the choice of the hyperparameters of the neural network and we observe qualitatively similar, but slightly inferior performance, when using a linear model instead of the neural network.
@@ 4.1 Imputation
Our analyses have established that LLMs have rich knowledge about entities such as regions and firms. Here we test whether this knowledge can be exploited to improve the imputation of missing values, a common challenge in economic modelling. As in Section 3, our approach does not assume the model has memorised the exact statistic to be imputed or that it can be recovered by prompting. Instead, we use the model’s embeddings as features, leveraging their generalised knowledge of entities. This makes the method applicable when the ground-truth statistic has not previously been produced or reported for certain entities and therefore can only be estimated.
To test whether embeddings can improve imputation accuracy, we compare two imputation strategies on our five datasets. The baseline strategy estimates imputed values using only the $K$ numeric features (e.g. population and unemployment rate) in the dataset. The embedding approach extends the feature matrix by appending the first $c$ PCA components of the embeddings.
To obtain embeddings for the entities, we use generic prompts (see Section 2.2), which only contain the name of the entity (e.g. region or firm) and the year. Although some of our datasets contain missing values, we do not impute them for evaluation because the true values are unknown, precluding accuracy assessment. Instead, we simulate missingness by randomly masking observed entries.
Our approach to introduce missing values is as follows: Given a dataset with $N$ observations and $K$ variables, all variables are transformed as described in Section 2.1 before they are normalised (mean of zero, standard deviation of 1). First, we sub-sample $k$ features, for each of which we randomly mask pp% of its values, creating the incomplete data $D_{miss}$. We then impute the missing values in $D_{miss}$ using imputation algorithm $m$.
We test three imputation methods $m$: using a low-rank Gaussian copula model (zhao2020matrix) and iterative multivariate imputation using (i) Bayesian Ridge regression and (ii) random forest (wilson2022miceforest). We assess the accuracy of the imputation with the mean absolute error (MAE) between the ground truth and the reconstructed values.
For each imputation method, we conduct a large array of experiments to ensure robustness of our results. Specifically, we consider all possible combinations of the parameter settings $k\in\{1,2,5\},p\in\{0.1,0.25,0.5\}$. To obtain stable results, we replicate the experiments for each parameter combination 25 times, randomly choosing the instances and features which are imputed. As embedding features, we use the first $c=25$ PCA components of the 25th layer of Llama 3 8B-Instruct.
Figure 8 presents the results across all five datasets and parameter settings $m$, $k$, and $p$. The embedding strategy consistently outperforms the baseline strategy. The improvement is particularly notable in some datasets. For instance, in the US counties dataset, the embedding strategy reduces the mean error by 14% on average across parameter settings.
@@ 4.2 Super-resolution
We define super-resolution as the estimation of statistics at a lower geographical level using information available at a higher level when data for the lower level are unavailable or unreliable. Like imputation, this task has high practical relevance. Many economic indicators are published only at higher administrative levels, despite systematic sub-regional heterogeneity that we may wish to estimate. LME, exploiting an LLM’s generalised knowledge about regions, could provide accurate estimates in this situation. For evaluation, however, lower-level ground truth is required, which is why we test our super-resolution approach only on data for which lower-level statistics are reported.
We approach the super-resolution task by training LME on a higher regional level (e.g. unemployment rate of US states) to obtain estimates at the lower regional level (e.g. unemployment rate of US counties). We use the completion prompt for both geographical levels. We test this approach on three of our datasets (US, EU, Germany) for which we were able to find data on two regional levels, as shown in Table 3.
@@
We compare this super-resolution approach with the text output at the lower level and with a naive method, which projects higher-level region values onto the lower level. For example, under the naive method, the unemployment rate of all counties in Alabama is set equal to the unemployment rate reported for the state of Alabama. Figure 9 presents the results, ordering the variables by decreasing performance of the text output.
The super-resolution approach performs well, surpassing the text output on most variables. A notable outlier is population, where the super-resolution performs worse than random. The naive approach performs well on many of the variables in Germany but falls behind the other two approaches on the US data. Given the small training sample size, in particular for the US and Germany (51 and 38 observations at the higher geographical level, respectively), the performance of the super-resolution approach is impressive. Super-resolution is a particular transfer learning approach and thus supports our previous finding (Section 3.4) that the embeddings can be used for accurate estimation outside the domain of the training set.
@@ 5 Conclusion
We have shown that LLMs know more than they say. A linear model trained on the LLM’s hidden states of only a few dozen labelled samples often outperforms the LLM’s text output in the estimation of statistics – in particular when the target variable is a less common statistic. A key contribution of our work is the transfer learning approach which shows that we can estimate target variables more accurately than the text output without any labelled data on that target variable. Finally, we provide robust evidence that the hidden states can be used for data processing tasks. In particular, adding a low-rank representation of the hidden states to the feature matrix consistently improves the performance of standard imputation approaches in all of our datasets. Our super-resolution results show that, in most cases, we can beat the text output in estimating statistics on a lower geographical level by learning from the higher geographical level.
We considered several LLMs with 1 to 70 billion parameters and show that our result that LLMs know more than they say holds across all of them, including a reasoning model that uses extensive test-time compute. We leave it for future work to test whether this finding also holds for the largest and most accurate open-source models such as Llama 3 405B or DeepSeek V3 with 671 billion parameters. In this study we exclusively worked with public data sources; it would be interesting to examine the extent to which our results hold for proprietary datasets, which are less likely to be observed in the training corpus of open-source LLMs. Finally, future research could investigate how useful the hidden states representing firms or geographic entities can be for other data processing tasks, such as outlier detection.
@@ A.1 US counties
The data on US counties for the year 2019 is one of our two main datasets and we used a variety of sources to assemble it. We obtained income and GDP data from the website of the Bureau of Economic Analysis [^10] https://www.bea.gov/data. The unemployment data was sourced from the Economic Research Service of the US Department of Agriculture [^11] https://www.ers.usda.gov/data-products/county-level-data-sets. Life expectancy data is provided by the Institute for Health Metrics and Evaluation in the Global Health Data Exchange database. We obtained mortgage delinquency rates from the Consumer Financial Protection Bureau. [^12] https://www.consumerfinance.gov/data-research/mortgage-performance-trends/download-the-data Data on the number of businesses was downloaded from the County Business Patterns Tables from the United States Census Bureau. [^13] https://www.census.gov/programs-surveys/cbp/data/tables.html Information on the median rent is available from the website of the Office of Policy Development and Research. [^14] https://www.huduser.gov/portal/datasets/50per.html The remaining variables we use come from the R package usdata (usdata) which lists the US Census and the American Community Survey as sources.